OpenStudy (anonymous):

1. (TCO 5 & 6) A cannon ball is dropped from a building. Suppose that the height of the cannon ball (in meters) after t seconds is given by the quadratic function f(t) = -4.4t^2 + 50 Find the average velocity in the time interval [1, 4] seconds. (Points : 5) -5.66 m/s -11.73 m/s -13.75 m/s -18.44 m/s -22 m/s

OpenStudy (anonymous):

\[v _{avg}=1/2(v _{f} + v _{i})\] \[v(t)=\frac {df}{dt}=-8.8t\]so the speed at t = 1 sec and t = 4 sec are\[v(1)=-8.8m/s^{2}(1s)=8.8m/s\]and\[v(4)=-8.8m/s^{2}(4s)=-35.2m/s\]plug into equation for v avg\[v _{avg}=1/2(-8.8m/s-35.2m/s)=22m/s\]